OpenStudy (vortish):

evaluate the function f(x)=9x+1 of the given values of the independent variables and simplify a) f(-5) b) f(x-5) c) f(-x)

OpenStudy (vortish):

i think I have a and b done a) would be -44 b) would be 9x+45

OpenStudy (anonymous):

A is correct! and B is sooo close! you forgot about the +1

OpenStudy (anonymous):

and that it's -5

OpenStudy (anonymous):

f(x-5) = 9(x-5)+1 = 9x -45+1 do the last step f(-x) = 9(-x) +1 = -9x+1